OpenStudy (anonymous):

Simplify the expression -1/8 - 2/7

OpenStudy (anonymous):

So you'll want to find a number they both have in common, the LCF, to be exact. to skip ahead, you'll find the 56 is the LCF. so \[\frac{ -1 }{8 } * \frac{ 7 }{ 7 }\] \[\frac{ -2 }{ 7 } * \frac{ 8 }{ 8 }\] for the first one, that would come out to -7/56 and the second; -16/56

OpenStudy (anonymous):

then just subtract -16 from -7 (the denominator [56] stays the same) so -23/56 is your answer, and that's as low as it can go